# Sobolev regularity theory for stochastic reaction-diffusion-advection equations with spatially homogeneous colored noises and variable-order nonlocal operators

### v2

@@ 1.4. Main results
 |($0,\tau\text{$]$\kern-1.49994pt$]$}:=\{(\omega,t):0<t\leq\tau(\omega)\},\quad\text{$[$\kern-3.80005pt$[$}0,\tau\text{$]$\kern-1.49994pt$]$}:=\{(\omega,t):0\leq t\leq\tau(\omega)\}.$
@@
Assumption 1.11.
There exist $c_{0}>0$ and $\delta_{0}\in(0,1]$ such that
$c_{0}\left(\frac{R}{r}\right)^{\delta_{0}}\leq\frac{\phi(R)}{\phi(r)},\quad\forall 0<r\leq R<\infty.$ (1.17)
@@
$\|f\|_{B_{p,q}^{\phi,\gamma-2/q}(\mathbb{R}^{d})}:=\|S_{0}f\|_{L_{p}(\mathbb{R}^{d})}+\left(\sum_{i=1}^{\infty}\phi(2^{2j})^{(q\gamma-2)/2}\|\Delta_{j}f\|_{L_{p}(\mathbb{R}^{d})}^{q}\right)^{1/q}<\infty.$
@@
where $\Psi$ is a function in $\mathcal{S}(\mathbb{R}^{d})$ which satisfies $supp(\mathcal{F}[\Psi])\subseteq\{\xi\in\mathbb{R}^{d}:2^{-1}\leq|\xi|\leq 2\}$ and
@@
Let $(F,B,\varphi)=(F(\omega,t,x,u),B(\omega,t,x,u),\varphi(\omega,t,x,u))$ be $\mathcal{P}\times\mathcal{B}(\mathbb{R}^{d})\times\mathcal{B}(\mathbb{R})$-measurable functions satisfying
@@
    $-F(u)\geq c_{\mathrm{s.d.}}|u|^{1+\lambda_{\mathrm{s.d.}}},\quad|B(u)|\leq c_{\mathrm{b.}}|u|^{1+\lambda_{\mathrm{b.}}},\quad|\varphi(u)|\leq c_{\mathrm{s.m.}}|u|^{1+\lambda_{\mathrm{s.m.}}},$
@@
    $\|b(\omega,t,\cdot)\|_{C^{2}(\mathbb{R}^{d})}+\||\zeta(\omega,t,\cdot)|+|\xi(\omega,t,\cdot)|\|_{L_{\infty}(\mathbb{R}^{d})}\leq K.$
@@
Let $p\in[2,\infty)$, $q\in(2,\infty)$, $\delta_{0}\in(0,1]$, $\gamma\in(0,1)$ and
@@
Suppose that Assumptions 1.5($\delta_{0},\gamma,r$), 1.11, 1.13 and 1.14 ($\delta_{0},\gamma$) hold. Suppose also that
@@
    $\frac{1}{q}\leq\alpha<\beta\leq\frac{\gamma}{2}-\frac{d}{2\delta_{0}p},$
@@
    $|u|_{C_{t,x}^{\alpha-\frac{1}{q},\phi_{\gamma-2\beta}-\frac{d}{p}}([0,\tau]\times\mathbb{R}^{d})}:=|u|_{C([0,\tau]\times\mathbb{R}^{d})}+\sup_{|h_{1}|\leq\tau,|h_{2}|\leq 1}\frac{|\mathcal{D}_{(h_{1},h_{2})}^{L}v|_{C([0,\tau]\times\mathbb{R}^{d})}}{|h_{1}|^{\alpha-\frac{1}{q}}+\phi(|h_{2}|^{-2})^{-\gamma+2\beta}|h_{2}|^{-\frac{d}{p}}},$
    $\mathcal{D}_{(h_{1},h_{2})}^{L}v(t,x):=\mathcal{D}_{(h_{1},h_{2})}(\mathcal{D}_{(h_{1},h_{2})}^{L-1}v)(t,x)$, $\mathcal{D}_{(h_{1},h_{2})}v(t,x):=v(t+h_{1},x+h_{2})-v(t,x)$ and $L$ is a natural number greater than $\gamma-d/p$. If $\alpha=1/q$, then
    $|u|_{C_{t,x}^{\alpha-\frac{1}{q},\phi_{\gamma-2\beta}-\frac{d}{p}}([0,\tau]\times\mathbb{R}^{d})}:=|u|_{C([0,\tau]\times\mathbb{R}^{d})}+\sup_{|h|\leq 1}\frac{|\mathcal{D}_{h}^{L}v|_{C([0,\tau]\times\mathbb{R}^{d})}}{\phi(|h|^{-2})^{-\gamma+2\beta}|h|^{-\frac{d}{p}}},$
    where $\mathcal{D}hv(t,x):=v(t,x+h)-v(t,x)$.
@@
We offer several comments on Theorem 1.15.
@@
    As demonstrated in Theorem 1.15, the strongly reinforced Dalang’s condition provides more flexibility in selecting the nonlinearity parameter $\lambda_{\mathrm{s.m.}}$. Specifically, from the choice of $\gamma$ and $r$, we derive a detailed characterization of noises $W$ so that we obtain the range of $\lambda_{\mathrm{s.m.}}$ depending on the noises and other parameters. Heuristically, if we take a properly narrow class (equivalently, $1<r$) of noises, we can obtain larger range of $\lambda_{\mathrm{s.m}}$.
@@
    Theorem 1.15 provides a more general framework compared to [9, Theorem 3.10] and [30, Theorem 3.5]. To facilitate a detailed comparison, we suggest an example case; set
    $d=2,\quad\gamma=\frac{1}{4},\quad\phi(\lambda)=\lambda,\quad\lambda_{\mathrm{b.}}=0,\quad p=q>16.$
    Suppose that spatially homogeneous colored noise $W$ with a nonnegative definite correlation $\pi$ satisfies
    $\int_{|x|<1}|x|^{-\frac{4}{3}}\mu(\mathrm{d}x)<\infty.$ (1.22)
    Here, $\mu$ denotes the measure associated with $\pi$.
    (Comparison with [9, Theorem 3.10]). We adopt the following parameter configuration;
    $\lambda_{\mathrm{s.m.}}=\frac{1}{8},\quad\lambda_{\mathrm{s.d.}}=0,\quad r=\frac{4}{3}.$
    These parameters simultaneously satisfy the conditions of [9, Theorem 3.10] and (1.19). Under this setup, the class of spatially homogeneous colored noise $W$ with a nonnegative definite correlation $\pi$ in [9, Theorem 3.10] meets the condition
    $\int_{|x|<1}|x|^{-\frac{5}{3}}\mu(\mathrm{d}x)<\infty.$
    To further clarify, if we assume $\mu(\mathrm{d}x):=|x|^{-\alpha}\mathrm{d}x$, the implications can be summarized as follows:
    [9, Theorem 3.10] $\implies 0<\alpha<\frac{1}{3},$
    Theorem 1.15 $\implies 0<\alpha<\frac{2}{3}.$
    (Comparison with [30, Theorem 3.5]). If the noise $W$ adheres to (1.22), it also fulfills [30, Assumption 2.3]. Assuming $\lambda_{\mathrm{s.d.}}=1$, the range of $\lambda_{\mathrm{s.m.}}$ allowed by [30, Theorem 3.5] is specified as
@@
  3. (iii)
    (Comparison with [27, Theorem 3.2]). Consider the stochastic reaction-diffusion equation given by
    $\partial_{t}u=\Delta^{\alpha/2}u+|u|^{1+\lambda_{\mathrm{s.m.}}}\dot{W},$
    where $\dot{W}$ represents the space-time white noise. It is noteworthy that [27, Theorem 3.2] excludes the case $\alpha=1$. In contrast, Theorem 1.15 explicitly accommodates this scenario, allowing $\alpha=1$ and thus extending the applicability of the results to a broader range of reaction-diffusion dynamics.
@@ 1.5. Strategy and Organization
We begin by outlining the proof concepts for Theorem 1.15. Our initial step involves constructing a local solution $u_{m}$ as detailed in Theorem 3.1-$(i)$, applying the established solvability of semilinear stochastic integro-differential equations in mixed norm $L_{q}(L_{p})$-spaces (Theorem 2.10). This solvability derives from the $H^{\infty}$-calculus discussed in Section 2.3. Following this, we establish the positivity of the local solution $u_{m}$ using a modified version of Krylov’s maximum principle for stochastic integro-differential equations (Theorem 3.1-($ii$)). The continuity of $u_{m}$ is then addressed in Theorem 3.1-($iii$). By leveraging the established positivity and continuity, we demonstrate that the explosion does not occur for the local solution $u_{m}$ (Theorem 3.1-($iv$)). Using the non-explosivity of the local solution, we proceed to derive the main results as formulated in Theorem 1.15.

### v3

@@ 1.4. Main results
[…]
@@
Remark 1.10.
The pair $(f,g)$ in (1.17) is uniquely determined by $u$. Indeed, if two pairs $(f_{1},g_{1})$ and $(f_{2},g_{2})$ satisfy (1.17), then for all $\varphi\in\mathcal{S}(\mathbb{R}^{d})$,
$\int_{0}^{t}(f_{2}(s,\cdot)-f_{1}(s,\cdot),\varphi)\,\mathrm{d}s=\sum_{k=1}^{\infty}\int_{0}^{t}(g_{1}^{k}(s,\cdot)-g_{2}^{k}(s,\cdot),\varphi)\,\mathrm{d}w_{s}^{k},$ (1.18)
for all $t\in[0,\tau]$ almost surely. Here, the left-hand side is a process of finite variation, while the right-hand side is a continuous local martingale. By [49, Proposition 1.2, Chapter IV], both processes must vanish identically. Thus $(f,g)$ is uniquely determined by $u$, justifying the notation $\mathbb{D}u$ and $\mathbb{S}u$.
@@
Assumption 1.12 ($\delta_{0}$).
The function $\phi:(0,\infty)\to[0,\infty)$ is a Bernstein function satisfying
$c_{0}\left(\frac{R}{r}\right)^{\delta_{0}}\leq\frac{\phi(R)}{\phi(r)},\quad\forall 0<r\leq R<\infty,\quad(c_{0}>0).$ (1.19)
This condition is a weak lower scaling condition for the Bernstein function $\phi$ with index $\delta_{0}\in(0,1]$. It ensures that $\phi$ grows at least polynomially at all scales and is a standard hypothesis in the study of nonlocal operators generated by subordinate Brownian motions.
Example 1.13.
The following Bernstein functions satisfy Assumption 1.12 $(\delta_{0})$. For additional examples and details, see [57, Chapter 16].
  1. (1)
    Stable subordinators: $\phi(\lambda)=\lambda^{\beta}$, $0<\beta\leq 1$ $\implies$ Assumption 1.12 $(\beta)$ holds.
  2. (2)
    Sum of stable subordinators: $\phi(\lambda)=\lambda^{\beta_{1}}+\lambda^{\beta_{2}}$, $0<\beta_{1},\beta_{2}\leq 1$ $\implies$ Assumption 1.12 $(\min(\beta_{1},\beta_{2}))$ holds.
  3. (3)
    Stable with logarithmic correction: $\phi(\lambda)=\lambda^{\beta}(\log(1+\lambda))^{\gamma}$, $\beta\in(0,1)$, $\gamma\in(-\beta,1-\beta)$ $\implies$ Assumption 1.12 $(\beta)$ holds.
  4. (4)
    Relativistic stable subordinators: $\phi(\lambda)=(\lambda+m^{1/\beta})^{\beta}-m$, $\beta\in(0,1)$, $m>0$ $\implies$ Assumption 1.12 $(\beta)$ holds.
  5. (5)
    Conjugate geometric stable subordinators: $\phi(\lambda)=\frac{\lambda}{\log(1+\lambda^{\beta/2})}$, $\beta\in(0,2)$ $\implies$ Assumption 1.12 $(1-\frac{\beta}{2})$ holds.
@@
$\|f\|_{B_{p,q}^{\phi,\gamma-2/q}(\mathbb{R}^{d})}:=\|S_{0}f\|_{L_{p}(\mathbb{R}^{d})}+\left(\sum_{j=1}^{\infty}\phi(2^{2j})^{(q\gamma-2)/2}\|\Delta_{j}f\|_{L_{p}(\mathbb{R}^{d})}^{q}\right)^{1/q}<\infty.$
@@
and $S_{0}:=\sum_{j\leq 0}\Delta_{j}$, where $\Psi$ is a function in $\mathcal{S}(\mathbb{R}^{d})$ which satisfies $supp(\mathcal{F}[\Psi])\subseteq\{\xi\in\mathbb{R}^{d}:2^{-1}\leq|\xi|\leq 2\}$ and
@@
Let $(F,B,\varphi)=(F(\omega,t,x,u),B(\omega,t,x,u),\varphi(\omega,t,x,u))$ be $\mathcal{P}\times\mathcal{B}(\mathbb{R}^{d})\times\mathcal{B}(\mathbb{R})$-measurable functions satisfying the followings:
@@
    $F(u)\geq c_{\mathrm{s.d.}}|u|^{1+\lambda_{\mathrm{s.d.}}},\quad|B(u)|\leq c_{\mathrm{b.}}^{1}|u|+c_{\mathrm{b.}}^{2}|u|^{1+\lambda_{\mathrm{b.}}},\quad|\varphi(u)|\leq c_{\mathrm{s.m.}}^{1}|u|+c_{\mathrm{s.m.}}^{2}|u|^{1+\lambda_{\mathrm{s.m.}}},$
@@
    $\|\vec{b}(\omega,t,\cdot)\|_{C^{2}(\mathbb{R}^{d};\mathbb{R}^{d})}+\||\zeta(\omega,t,\cdot)|+|\xi(\omega,t,\cdot)|\|_{L_{\infty}(\mathbb{R}^{d})}\leq K.$
@@
Let $p,q\in(2,\infty)$, $\delta_{0}\in(0,1]$, $\gamma\in(0,1)$ and
@@
Suppose that Assumptions 1.5 ($\delta_{0},\gamma,r$), 1.12 ($\delta_{0}$), 1.15 and 1.16 ($\delta_{0},\gamma$) hold. Suppose also that
@@
    $\frac{1}{q}\leq\alpha<\beta\leq\frac{\gamma}{2}-\frac{d}{2\delta_{0}p},$ (1.24)
@@
    $|u|_{C_{t,x}^{\alpha-\frac{1}{q},\phi_{\gamma-2\beta}-\frac{d}{p}}([0,\tau]\times\mathbb{R}^{d})}:=|u|_{C([0,\tau]\times\mathbb{R}^{d})}+[u]_{C_{t,x}^{\alpha-\frac{1}{q},\phi_{\gamma-2\beta}-\frac{d}{p}}([0,\tau]\times\mathbb{R}^{d})},$
    where
    $u]_{C_{t,x}^{\alpha-\frac{1}{q},\phi_{\gamma-2\beta}-\frac{d}{p}}([0,\tau]\times\mathbb{R}^{d})}:=\sup_{(t,x),(s,y)\in[0,\tau]\times\mathbb{R}^{d}}\frac{|u(t,x)-u(s,y)|}{|t-s|^{\alpha-\frac{1}{q}}+\phi(|x-y|^{-2})^{-\frac{\gamma}{2}+\beta}|x-y|^{-\frac{d}{p}}}.$
    If $\alpha=1/q$, then
    $|u|_{C_{t,x}^{\alpha-\frac{1}{q},\phi_{\gamma-2\beta}-\frac{d}{p}}([0,\tau]\times\mathbb{R}^{d})}:=|u|_{C([0,\tau]\times\mathbb{R}^{d})}+\sup_{t\in[0,\tau]}\sup_{x,y\in\mathbb{R}^{d}}\frac{|u(t,x)-u(t,y)|}{\phi(|x-y|^{-2})^{-\frac{\gamma}{2}+\beta}|x-y|^{-\frac{d}{p}}}.$
@@
We make several remarks on Theorem 1.17. The parameters appearing in (1.20)–(1.22) have the following interpretation:
  • $p,q$: integrability exponents for the solution $u$ (space and time, respectively);
  • $\gamma$: spatial regularity index of $u$;
  • $\delta_{0}$: weak lower scaling exponent of $\phi$;
  • $r$: spatial smoothness parameter of the noise $\dot{W}$ (via its correlation class);
  • $\lambda_{\mathrm{s.d.}},\,\lambda_{\mathrm{b.}},\,\lambda_{\mathrm{s.m.}}$: nonlinearity exponents of the strong dissipative term, the generalized Burgers-type term, and the superlinear multiplicative noise term, respectively.
@@
    (Effect of $\phi$ on $\lambda_{\mathrm{b.}}$ and $\lambda_{\mathrm{s.m.}}$) By the Mikhlin multiplier theorem and Assumption 1.12 $(\delta_{0})$, the Fourier multiplier
    $m(\xi):=\frac{(1+|\xi|^{2})^{\delta_{0}/2}}{(1+\phi(|\xi|^{2}))^{1/2}}$
    is bounded on $L_{p}(\mathbb{R}^{d})$ for $1<p<\infty$. Hence there exists $C=C(d,p,\delta_{0},\phi)$ such that
    $\|u\|_{L_{p}(\mathbb{R}^{d})}+\|\Delta^{\delta_{0}/2}u\|_{L_{p}(\mathbb{R}^{d})}\leq C\,(\|u\|_{L_{p}(\mathbb{R}^{d})}+\|\phi(\Delta)^{1/2}u\|_{L_{p}(\mathbb{R}^{d})}).$
    In particular, $\delta_{0}$ provides a baseline spatial regularity for every $u\in H_{p}^{\phi,1}(\mathbb{R}^{d})$. Thus, as $\delta_{0}\uparrow 1$ (i.e., as $\phi(\Delta)$ becomes more regularizing), the operator $\phi(\Delta)$ enforces stronger smoothing.
    Consequently, the constraint (1.20) becomes less restrictive for larger $\delta_{0}$, thereby enlarging the admissible range of the nonlinear exponents—most notably $\lambda_{\mathrm{b.}}$.
    Similarly, for the multiplicative noise constraint (1.21), note that the admissible window for $r$ contains
    $1\leq r\;<\;\frac{d}{\,d-\delta_{0}(1-\gamma)\,}.$
    Since the upper bound on $r$ increases with $\delta_{0}$, choosing $r$ close to this threshold makes (1.21) less restrictive as $\delta_{0}$ grows, which in turn enlarges the admissible range of $\lambda_{\mathrm{s.m.}}$.
@@
    (Effect of $r$ on $\lambda_{\mathrm{s.m.}}$) By the multiplicative-noise constraint (1.21), choosing smoother noise (i.e., larger $r$) makes the constraint less restrictive, thereby yielding a broader admissible range for $\lambda_{\mathrm{s.m.}}$ (with all other parameters fixed).
  3. (iii)
    The condition (1.22) guarantees the existence of parameters $\alpha,\beta$ such that (1.24) holds.
We conclude this subsection with a comparison of Theorem 1.17 to existing results. In particular, Theorem 1.17 provides a more general framework than the following:
[…]
For a concrete side-by-side comparison, we adopt the parameter choice
$d=2,\qquad\phi(\lambda)=\lambda^{\delta_{0}},\qquad p=q=16.$ (1.25)
1. Comparison with [9, Theorem 3.10].
Case 1 (fixed $\gamma$ and $\lambda_{\mathrm{s.m.}}$). We adopt the following parameter configuration:
$\delta_{0}=1,\qquad\gamma=\frac{1}{4},\qquad\lambda_{\mathrm{s.m.}}=\frac{1}{8},\qquad\lambda_{\mathrm{s.d.}}=0,\qquad\lambda_{\mathrm{b.}}=0,\qquad r=\frac{4}{3}.$ (1.26)
  • Assumption 1.5 $(\delta_{0},\gamma,r)$. Here
    $\alpha_{\delta_{0},\gamma,r}=d-r\big{(}d-\delta_{0}(1-\gamma)\big{)}=2-\frac{4}{3}\Big{(}2-\frac{3}{4}\Big{)}=\frac{1}{3}<\frac{d}{2}=1,$
    hence the correlation $\pi$ must satisfy
    $\int_{|x|<1}|x|^{-4/3}\,\pi(\mathrm{d}x)<\infty.$ (1.27)
  • Assumption in [9, Theorem 3.10]. To apply [9, Theorem 3.10], one requires
    $\int_{|x|<1}|x|^{-5/3}\,\pi(\mathrm{d}x)<\infty.$ (1.28)
Under (1.25), (1.26), and (1.27), all hypotheses of Theorem 1.17 are satisfied. Likewise, (1.25), (1.26), and (1.28) satisfy the assumptions of [9, Theorem 3.10].
If, moreover, $\pi(\mathrm{d}x)=|x|^{-\alpha}\,\mathrm{d}x$, then near the origin
$\int_{|x|<1}|x|^{-\beta}\,\pi(\mathrm{d}x)<\infty\quad\Longleftrightarrow\quad\beta+\alpha<d\ (=2).$
Consequently,
[…]
Thus, Theorem 1.17 applies to a strictly broader class of noises $\dot{W}$ than [9, Theorem 3.10].
Case 2 (fixed $\pi$ and $\lambda_{\mathrm{s.m.}}$). We adopt the following parameter configuration:
$\delta_{0}=1,\qquad\lambda_{\mathrm{s.m.}}=\frac{1}{8},\qquad\lambda_{\mathrm{s.d.}}=0,\qquad\lambda_{\mathrm{b.}}=0,\qquad r=1.$ (1.29)
Suppose $\pi(\mathrm{d}x)=|x|^{-1/4}\,\mathrm{d}x$ and $\gamma\in(0,1)$.
  • Condition (1.22). Here
    $\frac{2}{q}+\frac{d}{\delta_{0}p}=\frac{2}{16}+\frac{2}{16}=\frac{1}{4}<\gamma.$
  • Assumption in [9, Theorem 3.10]. One requires
    $\int_{|x|<1}|x|^{-\frac{4}{3}(1+\gamma)}\,\pi(\mathrm{d}x)<\infty\quad\Longleftrightarrow\quad\frac{4}{3}(1+\gamma)+\frac{1}{4}<2\ \Longleftrightarrow\ \gamma<\frac{5}{16},$
    and
    $\frac{d+2}{\gamma}=\frac{4}{\gamma}<p=16\quad\Longleftrightarrow\quad\gamma>\frac{1}{4}.$
Therefore, the admissible ranges of $\gamma$ are
[…]
Thus, Theorem 1.17 applies to a strictly broader range of regularity indices $\gamma$ than [9, Theorem 3.10].
2. Comparison with [30, Theorem 3.5].
Let us assume (1.25) and
$\delta_{0}=1,\qquad\gamma=\frac{1}{4},\qquad\lambda_{\mathrm{s.d.}}=1,\qquad\lambda_{\mathrm{b.}}=0,\qquad r=\frac{4}{3}.$
If the correlation $\pi(\mathrm{d}x):=|x|^{-1/3}\mathrm{d}x$, then $\pi$ satisfies (1.27) and [30, Assumption 2.3], simultaneously. Since $\lambda_{\mathrm{s.d.}}=1$, the range of $\lambda_{\mathrm{s.m.}}$ allowed by [30, Theorem 3.5] is
@@
3. Comparison with [27, Theorem 3.2].
Assume
$d=1,\qquad\phi(\lambda)=\lambda^{\delta_{0}},$ (1.30)
and take
$p=q=16,\qquad\delta_{0}=\frac{1}{2},\qquad\gamma=\frac{1}{2},\qquad\lambda_{\mathrm{s.m.}}=\frac{1}{32},\qquad\lambda_{\mathrm{b.}}=\lambda_{\mathrm{s.d.}}=0,\qquad r=1.$
Let $\pi(\mathrm{d}x)=\varepsilon_{0}(\mathrm{d}x)$ denote the centered Dirac measure. Then all assumptions of Theorem 1.17 are satisfied, so the theorem applies in this setting. By contrast, [27, Theorem 3.2] excludes the case $\delta_{0}=\frac{1}{2}$.
Next, still under (1.30) with $\pi(\mathrm{d}x)=\varepsilon_{0}(\mathrm{d}x)$, consider
$p=q=64,\qquad\delta_{0}=\frac{2}{3},\qquad\gamma=\frac{1}{12},\qquad\lambda_{\mathrm{b.}}=\lambda_{\mathrm{s.d.}}=0,\qquad r=\frac{4}{3}.$
In this case, the admissible range for $\lambda_{\mathrm{s.m.}}$ under Theorem 1.17 is
$0<\lambda_{\mathrm{s.m.}}<\frac{9}{64},$
whereas the corresponding range under [27, Theorem 3.2] is
$0<\lambda_{\mathrm{s.m.}}<\frac{1}{12}.$
Since $\frac{9}{64}>\frac{1}{12}$, Theorem 1.17 permits strictly larger superlinear growth in the multiplicative noise term than [27, Theorem 3.2].
@@ 1.5. Strategy and Organization
We begin by outlining the proof concepts for Theorem 1.17. Our initial step involves constructing a local solution $u_{m}$ as detailed in Theorem 3.1-$(i)$, applying the established solvability of semilinear stochastic integro-differential equations in mixed norm $L_{q}(L_{p})$-spaces (Theorem 2.10). This solvability derives from the $H^{\infty}$-calculus discussed in Section 2.3. Following this, we establish the nonnegativity of the local solution $u_{m}$ using a modified version of Krylov’s maximum principle for stochastic integro-differential equations (Theorem 3.1-($ii$)). The continuity of $u_{m}$ is then addressed in Theorem 3.1-(iiiiii). By leveraging the established nonnegativity and continuity, we demonstrate that the explosion does not occur for the local solution $u_{m}$ (Theorem 3.1-($iv$)). Using the non-explosivity of the local solution, we proceed to derive the main results as formulated in Theorem 1.17.